all the cube root of iota (i) in a complex number (C) are z1= cosπ/2 + i sinπ/2 , z2= cosπ/6 +i sinπ/6 and z3= cos5π/6 + i sin5π/6 .


Verify if the statement is true or false. Give reason for your answer in the form of a short proof or a counterexample.

1
Expert's answer
2021-11-08T19:44:19-0500

Solution:

We know that

"i^{1/3}=\cos(4k+1)\dfrac{\pi}6+i\sin(4k+1)\dfrac{\pi}6; k=0,1,2"

So, for k=0

"i^{1/3}=\cos\dfrac{\pi}6+i\sin \dfrac{\pi}6
\\=\dfrac{\sqrt3}2+\dfrac i2"

For k=1

"i^{1/3}=\cos\dfrac{5\pi}6+i\sin \dfrac{5\pi}6
\\=-\cos\dfrac{\pi}6+i\sin \dfrac{\pi}6
\\=-\dfrac{\sqrt3}2+\dfrac i2"

For k=2

"i^{1/3}=\cos\dfrac{9\pi}6+i\sin \dfrac{9\pi}6
\\=\cos\dfrac{3\pi}2+i\sin \dfrac{3\pi}2
\\=-\cos\dfrac{\pi}2+i(-\sin \dfrac{\pi}2) 
\\=-i"

Given ones are: z1= cosπ/2 + i sinπ/2 , z2= cosπ/6 +i sinπ/6 and z3= cos5π/6 + i sin5π/6 .

z1 is incorrect, rest are correct.

So, it is false.